OpenStudy (anonymous):

How many molecules (not moles) of NH3 are produced from 3.10×10−4g of H2?

OpenStudy (cuanchi):

\[3.10 \times10^{-4}g H _{2}\times \frac{ 2 mol H atom }{ 2g H _{2} }\times \frac{ 1mol NH _{3} }{ 3mol H atom }\times \] \[\frac{ 6.02\times 10^{23}molec NH _{3}}{ 1mol NH _{3} }=6.22\times10^{19}molecules of NH _{3} \]